OpenStudy (anonymous):

8-10). Perform the following multiplication. The answer should be in simplified form for radical expressions. sqrt(6) sqrt(7)

OpenStudy (anonymous):

sqrt(6)sqrt(7)=sqrt(6*7)=sqrt(42)